he heights of adult men in America are normally distributed, with a mean of 69.2 inches and a standard deviation of 2.65 inches. The heights of adult women in America are also normally -distributed, but with a mean of 64.3 inches and a standard deviation of 2.59 inches.
a) If a man is 6 feet 3 inches tall, what is his z-score (to two decimal places)?
z =

b) If a woman is 5 feet 11 inches tall, what is her z-score (to two decimal places)?
z =

c) Who is relatively taller?
-The 6 foot 3 inch American man
-The 5 foot 11 inch American woman

Answer by ewatrrr(24785)

Note: standard Normal Curve: µ = 0 and σ =1
for ex: to the right of center, z-score becomes larger as distance, in terms of σ from the mean, increases.

Men Normal Distribution: µ = 69.2 and σ = 2.65
Women Normal Distribution: µ = 64.3 and σ = 2.59
z(man 75") = ( 75 - 69.2)/2.65 = 2.1887
z(woman 71") = (71 - 64.3)/2.59 = 2.5869****Represents More standard deviations from mean
c) Who is 'relatively taller?
-The 5 foot 11 inch American woman